how do i graph the following 2 equations?
equation 1. y = 4x + 3
equation 2. y + x = -5

Answer by Cromlix(4381)

By placing consecutive values of x into the equations
and generate y values.
Example 1 using x = -2 => y = 4x + 3 => y = 4(-2) + 3 = -5
using x = -1 => y = 4x + 3 => y = 4(-1) + 3 = -1
using x = 0 => y = 4x + 3 => y = 4(0) + 3 = 3
using x = 1 => y = 4x + 3 => y = 4(1) + 3 = 7
using x = 2 => y = 4x + 3 => y = 4(2) + 3 = 11
Now plot the points (-2,-5), (-1,-1), (0,3), (1,7) and (2,11)
on your graph.
Example 2. Sort equation into y = -x -5 (It must be arranged
to the y = mx + c form. m = gradient. c = where line
cuts the y axis)
using x = -2 => y = -x -5 => y = -(-2) -5 = -3
using x = -1 => y = -x -5 => y = -(-1) -5 = -4
using x = 0 => y = -x -5 => y = -(0) -5 = -5
using x = 1 => y = -x -5 => y = -(1) -5 = -6
using x = 2 => y = -x -5 => y = -(2) -5 = -7
Now plot the points (-2,-3), (-1,-4),(0,-5), (1,-6) and (2,-7)
on your graph.